OpenStudy (anonymous):

help me pleaseeeeeeeee;( 15. Solve 4y + 16x = 20 for y. Put your answer in the form y=mx+b with no spaces. y=-4x+5 y=-4x-5 y=4x-5 y=4x+5 16. Solve 2y = 4x - 24 for y. Put your answer in the form y=mx+b with no spaces. y=2x-12 y=-2x-12 y=2x+12 y=-2x+12 17. Solve 6x = 2y - 18 for y. Put your answer in the form y=mx+b with no spaces. y=-3x+9 y=3x+9 y=-3x-9 y=3x-9 18. What is the slope of this equation? y = 8x - 7 -7 -8 7 8

OpenStudy (anonymous):

#15 is y=5-4x #16 is y=2x-12 #17 is y=3x+9 #18 is 8

OpenStudy (anonymous):

thank you :*